# Efficient Crystal Structure Prediction Using Genetic Algorithm and Universal Neural Network Potential

### v1

@@ I Introduction
While CSP plays an integral role in predicting stable crystal structures, it represents a daunting global optimization task owning to the vast energy landscape. Substituting elements from known crystal structure prototypes is a prevailing method for generating candidate structures Hautier et al. (2011). Although the prototype substitution method often gives reasonable results, its coverage falls short, especially in multicomponent systems. Metaheuristic algorithms offer another approach to creating novel crystal structures. These include methods such as random structure search Pickard and Needs (2011), basin hopping Wales and Doye (1997), minima hopping Amsler and Goedecker (2010); Krummenacher et al. (2024), genetic algorithm (GA) Deaven and Ho (1995); Oganov and Glass (2006); Glass et al. (2006); Lonie and Zurek (2011); Falls et al. (2021), particle swarm optimization Wang et al. (2010), and Bayesian optimization Yamashita et al. (2018); Bisbo and Hammer (2020). Typically, these CSP methods are combined with density functional theory (DFT) calculations to evaluate the formation energy of the candidate structures. However, the time-consuming DFT calculations, which limit the exploration of structures, significantly hinder the efficiency of CSP.
@@
Recent advancements in machine learning potential, interatomic potential fitted by DFT calculations, enable an efficient approach to CSP thanks to their fast and accurate energy evaluation Behler and Parrinello (2007); Bartók et al. (2010); Shapeev (2016); Drautz (2019); Podryabinkin et al. (2019). Because we are often interested in multicomponent systems like ternary or quaternary systems in CSP, it is preferable for the machine learning potential to exhibit scalability with the number of elements and transferability across various systems. Also, high accuracy is required to capture subtle energy differences among distinct structures in CSP. A universal neural network potential (NNP) trained with extensive datasets has been gaining significant attention owing to its scalability and transferability, which includes M3GNet Chen et al. (2019); Chen and Ong (2022), ALIGNN-FF Choudhary et al. (2023), CHGNet Deng et al. (2023), MACE-MP-0 Batatia et al. (2022, 2023), GNoME Merchant et al. (2023), Orb Neumann et al. (2024), EquiformerV2-OMat24 Barroso-Luque et al. (2024); Schmidt et al. (2023), MatterSim Yang et al. (2024), and PFP Takamoto et al. (2022a, b); Jacobs et al. (2025).
@@
Applying a universal NNP to CSP is straightforward. For example, GN-OA Cheng et al. (2022) combines the existing metaheuristic algorithms with a universal NNP. GNoME Merchant et al. (2023) has unveiled various previously unreported stable structures with an extension of prototype substitution and random structure generation. Nevertheless, further development of CSP algorithms employing universal NNPs have room for enhancement. While the existing ab initio CSP methods often implicitly assume the number of energy evaluations to be limited due to the computational cost of DFT, the universal NNPs can allow us the vast number of energy evaluations and consider an entire convex hull concurrently, thereby appreciating the diversity of crystal structures all at once.
@@ III.2 Population filtering
In this step, we selected structures close to the convex hull with the most recently updated composition. Our algorithm design is based on two intuitions: (1) lower-energy structures should exist near the structures close to the convex hull, and (2) there should be some room for improvement in the neighbors of the compositions that observed recent updates on their lowest energy. Population filtering is similar to the aging evolution strategy Real et al. (2019) in that it ignores the structures that have not been updated for a long time.
@@ III.4 Crossover and mutation
Crossover and mutation are the processes of generating new structures $P_{n+1}$ from the elite population $Q_{n+1}$. Crossover involves combining two parent structures to generate a child structure, and mutation involves changing the structure of a parent structure.
The crossover and mutations used in our proposed method are listed in Table 2. Several methods are implemented in ASE Larsen et al. (2017), and others are our original methods. To customize the methods in ASE, we modify the implementation and release them as a package.  [^1] The modified crossover and mutation methods are available at https://pypi.org/project/pfn-ase-extras/
@@
at their unit cells along a crystal plane and joining them Vilhelmsen and Hammer (2012).
Modified ASE $1/2$
Random composition mutation Generate a child structure by randomly changing its composition. ASE $1/12$
@@
to the basis vectors.
ASE $1/12$
Rattle mutation Perturb the atomic positions by a random displacement. ASE $1/12$
Permutation mutation Swap the atomic positions of two selected atoms. ASE $1/12$
@@
Original $1/12$
Random structure generation mutation Generate a structure by the random structure generation in Section III.1. Original $1/12$
The cut-and-splice crossover is a method to generate a child structure by cutting two parent structures at their unit cells along a crystal plane and joining them. Originally, its implementation in ASE only allowed crossover between parent structures with the same composition. We modify the implementation to allow crossover between parent structures with different compositions.
@@ IV Results and discussion
In this section, we present the results of our proposed method for CSP using PFP and GA. First, we evaluated the performance of the proposed method by comparing it with a random search in Section IV.1. We then demonstrated the effectiveness of the proposed method by performing CSP searches for binary and ternary systems with MP in Section IV.2. The results demonstrate that the proposed method can efficiently explore the search space and discover new crystal structure candidates that update the convex hull of MP. For the energy evaluation, we used the energy corrections described in Section II.3.
IV.1 Quantitative comparison with random search and MP
Figure 4: Comparison between random search (orange) and the proposed method (blue) by the normalized hull volume history. The hatchings indicate the minimum and maximum normalized hull volumes within the five searches. The black dashed line indicates the normalized volumes for the convex hull constructed from MP structures.
@@ IV.1 Quantitative comparison with random search and MP
Figure 5: Distribution of the energy above the hull of the crystal structures of MP. The convex hull is reconstructed using the CSP results obtained using the method proposed in Section IV.1. The results of the five searches are aggregated and presented by the system.
@@
IV.2 CSP for binary and ternary systems
We demonstrated the capability of our CSP method using PFP to discover new crystal structure candidates for binary and ternary systems. The search conditions were the same as those in Section IV.1; however, the structures from MP are added to the initial population of the search. The crystal structures obtained from the search were re-evaluated using DFT calculations to verify their validity. We report the results for systems In–Li, As–V, Al–Li–Pd, and La–Mo–O shown in Figs. 6 and 7.
@@ IV.2 CSP for binary and ternary systems
Crystal structures that update the convex hull of MP in the systems shown in Figs. 6 and 7 are listed in Table 4 222 These crystal structures in xyz formats are available at Supplemental Materials. . The space-group type was identified using spglib with symprec = 0.01 and angle_tolerance = 5 Togo et al. (2024). Additionally, we matched these crystal structures with AFLOW prototypes Mehl et al. (2017); Hicks et al. (2019, 2021); Eckert et al. (2024) by assigning the AFLOW label using aviary.wren Goodall et al. (2022). Several crystal structures were not found as prototype structures in the AFLOW database, indicating that our CSP method can discover novel crystal structures. Some of the listed crystal structures are visualized in Fig. 8.

### v2

@@ I Introduction
While CSP plays an integral role in predicting stable crystal structures, it represents a daunting global optimization task owning to the vast energy landscape. Substituting elements from known crystal structure prototypes is a prevailing method for generating candidate structures Hautier et al. (2011). Although the prototype substitution method often gives reasonable results, its coverage falls short, especially in multicomponent systems. Metaheuristic algorithms offer another approach to creating novel crystal structures. These include methods such as random structure search Pickard and Needs (2011), basin hopping Wales and Doye (1997), minima hopping Amsler and Goedecker (2010); Krummenacher et al. (2024), genetic algorithm (GA) Deaven and Ho (1995); Oganov and Glass (2006); Glass et al. (2006); Oganov et al. (2011); Lyakhov et al. (2013); Lonie and Zurek (2011); Falls et al. (2021), particle swarm optimization Wang et al. (2010), and Bayesian optimization Yamashita et al. (2018); Bisbo and Hammer (2020). Typically, these CSP methods are combined with density functional theory (DFT) calculations to evaluate the formation energy of the candidate structures. However, the time-consuming DFT calculations, which limit the exploration of structures, significantly hinder the efficiency of CSP.
@@
Recent advancements in machine learning potential, interatomic potential fitted by DFT calculations, enable an efficient approach to CSP thanks to their fast and accurate energy evaluation Behler and Parrinello (2007); Bartók et al. (2010); Shapeev (2016); Drautz (2019); Podryabinkin et al. (2019). Because we are often interested in multicomponent systems like ternary or quaternary systems in CSP, it is preferable for the machine learning potential to exhibit scalability with the number of elements and transferability across various systems. Also, high accuracy is required to capture subtle energy differences among distinct structures in CSP. A universal neural network potential (NNP) trained with extensive datasets has been gaining significant attention owing to its scalability and transferability, which includes M3GNet Chen et al. (2019); Chen and Ong (2022), ALIGNN-FF Choudhary et al. (2023), CHGNet Deng et al. (2023), MACE-MP-0 Batatia et al. (2022, 2023), GNoME Merchant et al. (2023), Orb Neumann et al. (2024), EquiformerV2-OMat24 Barroso-Luque et al. (2024); Schmidt et al. (2023a), MatterSim Yang et al. (2024), and PFP Takamoto et al. (2022a, b); Jacobs et al. (2025).
@@
Applying a machine learning potential to CSP is relatively straightforward and it still achieves a significant improvement in efficiency Jacobsen et al. (2018); Podryabinkin et al. (2019); Tong et al. (2018); Wang et al. (2020); Park and Wolverton (2020); Kang et al. (2022); Cheng et al. (2022); Schmidt et al. (2023b); Merchant et al. (2023); Chang et al. (2024). Nevertheless, further development of CSP algorithms employing universal NNPs have room for enhancement. While the existing ab initio CSP methods often implicitly assume the number of energy evaluations to be limited due to the computational cost of DFT, the universal NNPs can allow us the vast number of energy evaluations and consider an entire convex hull concurrently, thereby appreciating the diversity of crystal structures all at once.
@@
In recent years, research on multi-objective genetic algorithms has been emerging Omee et al. (2023, 2025); Donaldson et al. (2024). Donaldson et al. Donaldson et al. (2024) proposed a method that combines a multi-objective genetic algorithm with a universal NNP to explore the entire convex hull of a given system. Exploring the entire convex hull requires significantly longer simulations than composition-fixed searches. However, the existing method quickly fall into local optima and are not able to benefit from prolonged simulations as shown in Section IV. We propose a method that allows long-time search without being trapped in local minima by dynamically shifting the search region via the forgetting of outdated structures.
Moreover, even with relatively long simulations, the search space of crystal structure prediction remains vast. Exhaustive search is of course infeasible, and random search is insufficient in terms of sample efficiency. To explore efficiently, it is essential to go beyond conventional genetic algorithms and exploit similarities between structures with different compositions. We propose a niching strategy that considers both proximity to the convex hull and similarities between formation energy and composition. While the idea of emphasizing structures near the convex hull is shared with Donaldson et al. Donaldson et al. (2024), our formulation was developed independently.
@@ III.2 Population filtering
In this step, we selected structures close to the convex hull with the most recently updated composition. Our algorithm design is based on two intuitions: (1) lower-energy structures should exist near the structures close to the convex hull, and (2) there should be some room for improvement in the neighbors of the compositions that observed recent updates on their lowest energy. Population filtering is similar to the aging evolution strategy Zhu et al. (2012); Real et al. (2019) in that it ignores the structures that have not been updated for a long time.
@@ III.4 Crossover and mutation
at their unit cells along a crystal plane and joining them Oganov et al. (2010, 2011).
$1/2$
Random composition mutation Generate a child structure by randomly changing its composition Glass et al. (2006); Oganov et al. (2010); Lyakhov et al. (2013). $1/12$
@@
to the basis vectors Glass et al. (2006); Oganov et al. (2010); Lyakhov et al. (2013).
$1/12$
Rattle mutation Perturb the atomic positions by a random displacement Glass et al. (2006); Oganov et al. (2010); Lyakhov et al. (2013). $1/12$
Permutation mutation Swap the atomic positions of two selected atoms Glass et al. (2006); Oganov et al. (2010); Lyakhov et al. (2013). $1/12$
@@
$1/12$
Random structure generation mutation Generate a structure by the random structure generation in Section III.1. $1/12$
Crossover and mutation are the processes of generating new structures $P_{n+1}$ from the elite population $Q_{n+1}$. Crossover involves combining two parent structures to generate a child structure, and mutation involves changing the structure of a parent structure. The crossover and mutations used in our proposed method are listed in Table 2. Several methods are implemented in ASE Larsen et al. (2017) and we modify some of the implementation 111 The modified crossover and mutation methods are released and available at https://pypi.org/project/pfn-ase-extras/. .
The cut-and-splice crossover is a method to generate a child structure by cutting two parent structures at their unit cells along a crystal plane and joining them Deaven and Ho (1995); Vilhelmsen and Hammer (2012). Because its implementation in ASE only allowed crossover between parent structures with the same composition, we modify it to allow crossover between parent structures with different compositions, which is already known in the field of CSP Oganov et al. (2010, 2011).
@@ IV Results and discussion
In this section, we present the results of our proposed method for CSP using PFP and GA. First, we conducted an ablation study of the population filtering in Section IV.1 to demonstrate the effectiveness of the proposed method. Secondly, we evaluated the performance of the proposed method by comparing it with a random search in Section IV.2. We then demonstrated the effectiveness of the proposed method by performing CSP searches for binary and ternary systems with MP in Section IV.3. The results demonstrate that the proposed method can efficiently explore the search space and discover new crystal structure candidates that update the convex hull of MP. For the energy evaluation, we used the energy corrections described in Section II.3.
IV.1 Ablation study of population filtering
Figure 4: Compositional dispersion history of the Br–Pb–Rh system. The Average Nearest Neighbor Index (ANNI) is plotted against the generation number. The blue line indicates the case with population filtering, while the purple line indicates the case without population filtering. The shaded area represents the standard deviation of ANNI values across 10 experiments. Lower ANNI values indicate greater compositional skewness.
Figure 5: Compositional distribution of structures explored at each generation in the Br–Pb–Rh system. The upper panel shows the case with population filtering, while the lower panel shows the case without population filtering. From left to right, the generations are 5, 25, 50, and 100.
In this section, we conducted an ablation study of the population filtering. Population filtering is performed immediately prior to elite population selection in each generation to preserve the diversity of candidate compositions. The ablation study compared two cases: one with population filtering and one without. We demonstrated that implementing population filtering enables exploration of a wider range of structures across various compositions compared to its absence.
In this experiment, we examined the Br–Pb–Rh system with the following parameters: maximum number of atoms in the unit cell was 128, the population size was 250, the total number of trials was 50,000, $\alpha=2.4$, and $\epsilon=0.25$. We repeated the search 10 times for each case to account for the stochastic nature of the CSP search. The genetic operations and their selection probabilities are specified in Table 2. We conducted comparisons under two conditions: with and without enabling population filtering as described in Section III.2.
To quantitatively assess how compositional dispersion evolves across generations, we employed the Average Nearest Neighbor Index (ANNI) Clark and Evans (1954) to evaluate the compositional homogeneity of structures explored in each generation. Lower ANNI values indicate greater compositional skewness. As shown in Figure 4, when population filtering is enabled, the ANNI values are significantly higher during early generations compared to when filtering is disabled, indicating greater compositional dispersion. However, beginning approximately from the 40th generation onward, ANNI values become comparable regardless of whether population filtering is active or not, suggesting that compositional distribution becomes similarly skewed throughout both search phases.
Figure 5 illustrates the distribution of compositions of structures explored at each generation for one search. At generations 5 and 25, we observe that enabling population filtering results in a more broadly distributed range of structural compositions. From generation 50 onward, even when population filtering is active, the distribution of compositions becomes narrower, exhibiting patterns similar to when population filtering is disabled. Conversely, when population filtering is disabled, the explored structures show concentration within a narrow compositional range, and this range remains relatively stable across generations. This demonstrates that population filtering enables exploration of structures across a wider portion of the search space.
IV.2 Quantitative comparison with random search and MP
Figure 6: Comparison between random search (blue) and the proposed method (orange) by the normalized hull volume history. The hatchings indicate the minimum and maximum normalized hull volumes within the five searches. The black dashed line indicates the normalized volumes for the convex hull constructed from MP structures.
@@ IV.2 Quantitative comparison with random search and MP
Figure 7: Distribution of the energy above the hull of the crystal structures of MP. The convex hull is reconstructed using the CSP results obtained using the method proposed in Section IV.2. The results of the five searches are aggregated and presented by the system.
@@
IV.3 CSP for binary and ternary systems
We demonstrated the capability of our CSP method using PFP to discover new crystal structure candidates for binary and ternary systems. The search conditions were the same as those in Section IV.2; however, the structures from MP are added to the initial population of the search. The crystal structures obtained from the search were re-evaluated using DFT calculations to verify their validity. We report the results for systems In–Li, As–V, Al–Li–Pd, and La–Mo–O shown in Figs. 8 and 9.
@@ IV.3 CSP for binary and ternary systems
Crystal structures that update the convex hull of MP in the systems shown in Figs. 8 and 9 are listed in Table 4 222 These crystal structures in xyz formats are available at Supplemental Materials. . The space-group type was identified using spglib with symprec = 0.01 and angle_tolerance = 5 Togo et al. (2024). Additionally, we matched these crystal structures with AFLOW prototypes Mehl et al. (2017); Hicks et al. (2019, 2021); Eckert et al. (2024) by assigning the AFLOW label using aviary.wren Goodall et al. (2022). Several crystal structures were not found as prototype structures in the AFLOW database, indicating that our CSP method can discover novel crystal structures. Some of the listed crystal structures are visualized in Fig. 10.